How much water needs to be added to a half ounce of a solution in which the active ingredient is 10% so as to make that half ounce of solution into having .5% of the active ingredient?
Answer by stanbon(75887)

How much water needs to be added to a half ounce of a solution in which the active ingredient is 10% so as to make that half ounce of solution into having .5% of the active ingredient?
-----
Equation:
active + active = active
0.10*(1/2) + 0*x = 0.005((1/2)+x))
-----------------------------------------
Multiply thru by 1000 to get:
50 + 0 = (5/2) + 5x
5x = 47.5
x = 9.5 ounces (amt of water needed)
-----------------------
Cheers,
Stan H.